# Forecasts for Galaxy Formation and Dark Matter Constraints from Dwarf Galaxy Surveys

### v1

@@ 9 Summary
We have forecasted the galaxy formation and DM constraints that future surveys of dwarf galaxy populations around MW-mass hosts can deliver. We find that:
@@ Acknowledgements
We thank Keith Bechtol for comments on the manuscript and Rui An, Shany Danieli, Alex Drlica-Wagner, Ivan Esteban, and Josh Simon for helpful discussions. VG acknowledges support from NASA through the Astrophysics Theory Program, Award Number 21-ATP21-0135 and from the National Science Foundation (NSF) CAREER Grant No. PHY-2239205. The work of LAM was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA. This work received additional support from the U.S. Department of Energy under contract number DE-AC02-76SF00515 to SLAC National Accelerator Laboratory and from the Kavli Institute for Particle Astrophysics and Cosmology at Stanford University. This work was performed in part at the Aspen Center for Physics, which is supported by NSF grant PHY-2210452.
The computations presented here were conducted through Carnegie’s partnership with the Resnick High Performance Computing Center, a facility supported by the Resnick Sustainability Institute at the California Institute of Technology. This work used data from the Symphony suite of simulations (http://web.stanford.edu/group/gfc/symphony/), which was supported by the Kavli Institute for Particle Astrophysics and Cosmology at Stanford University and SLAC National Accelerator Laboratory, and by the U.S. Department of Energy under contract number DE-AC02-76SF00515 to SLAC National Accelerator Laboratory.
@@ Appendix C Posterior Distributions
Finally, we present the full posteriors from each forecast described in the main text. For all posteriors below, dark (light) contours represent $68\%$ ($95\%$) confidence intervals, and dashed black lines show the input value of each parameter; contours are colored blue (red) for our one-host (two-host) forecasts. Note that $\sigma_{M}$ and $\sigma_{\log R}$ are reported in $\rm{dex}$, $\mathcal{M}_{50}$ is reported as $\log(\mathcal{M}_{50}/M_{\rm{\odot}})$, $\mathcal{A}$ is reported in $\rm{pc}$, $M_{\mathrm{hm}}$ is reported as $\log(M_{\mathrm{hm}}/M_{\rm{\odot}})$, and the remaining galaxy–halo connection and beyond-CDM parameters are dimensionless. Furthermore, the two-host posteriors below always combine hosts with different bright satellite populations (i.e., classical satellite abundances that differ at a level comparable to the $1\sigma$ host-to-host scatter in our simulation suite).
@@ C.1 Galaxy Formation Cutoff
The posteriors from our one and two-host “strong” galaxy formation cutoff forecasts are shown in Figures 10 and 11, respectively. The parameters of interest for this forecast (i.e., $\sigma_{M}$, $\mathcal{M}_{50}$, and $\mathcal{S}_{\mathrm{gal}}$) are not strongly correlated with the other galaxy–halo connection parameters; however, there is a prominent degeneracy between $\mathcal{M}_{50}$ and $\sigma_{M}$ (and, to a lesser extent, between $\mathcal{S}_{\mathrm{gal}}$ and $\sigma_{M}$). As described in Section 4, this follows because larger values of $\sigma_{M}$ preferentially cause faint galaxies to up-scatter to observable luminosities, decreasing the average masses of halos inferred to host the faintest observed galaxies (a form of Eddington bias). Increasing $\mathcal{M}_{50}$ cuts off galaxy formation more drastically in lower-mass halos, which is not allowed in regions of parameter space where the galaxies they host are crucial to explain the data, resulting in a negatively-sloped degeneracy between $\mathcal{M}_{50}$ and $\sigma_{M}$.
Next, the posteriors from our one and two-host “weak” galaxy formation cutoff forecasts are shown in Figures 12 and 13, respectively. In the one-host case, we only recover an upper limit on $\mathcal{M}_{50}$, and we observe weaker degeneracies between $\mathcal{M}_{50}$ and the remaining galaxy–halo connection parameters. Although uncertainties significantly shrink in the two-host forecast, $\mathcal{M}_{50}$ is still not constrained from below in this case, consistent with our discussion in Section 4. Note that the upper limit on $\mathcal{M}_{50}$ in the two-host forecast is not in significant tension with its input value.
@@
Free Parameter Prior Distribution Motivation
@@
WDM half-mode mass $\log(M_{\mathrm{hm}}/M_{\rm{\odot}})\sim\rm{unif}(5,10)$ Brackets current constraints (Banik et al., 2021b; Gilman et al., 2020a)
@@ C.2 Warm Dark Matter
The posteriors from our one and two-host WDM forecasts assuming a “weak” galaxy formation cutoff and CDM subhalo abundances are shown in Figures 14 and 15, and the corresponding “strong” galaxy formation cutoff runs are shown in Figures 16 and 17, respectively. The posteriors are qualitatively similar in these scenarios, and our description below applies to both.
@@
All of the same degeneracies among galaxy–halo connection parameters noted in Appendix C.1 remain when $M_{\mathrm{hm}}$ is added to the inference. In addition to the significant degeneracies between $M_{\mathrm{hm}}$ and both $\mathcal{M}_{50}$ and $\sigma_{M}$ discussed in Section 5, these posteriors also illustrate a weak degeneracy between $M_{\mathrm{hm}}$ and $\mathcal{S}_{\mathrm{gal}}$ that persists even when two complete satellite populations with different bright satellite populations are used. Figures 16 and 17 clearly illustrate that the $\mathcal{M}_{50}$$\sigma_{M}$ degeneracy is broken when combining two complete satellite population with differing classical satellite abundances, which in turn yields a much stronger upper limit on $M_{\mathrm{hm}}$ in the two-host case.
@@
Figures 18 and 19 respectively show the posteriors from our one and two-host WDM forecasts assuming a “strong” galaxy formation cutoff and WDM subhalo abundances with $M_{\mathrm{hm}}=10^{8}~{}M_{\mathrm{\odot}}$ ($m_{\mathrm{WDM}}=4.9~{}\mathrm{keV}$). The degeneracies described in the previous paragraph all remain in this case and often become more prominent. As discussed in Section 5, this follows because, in our model, astrophysical and DM-induced cutoffs affect dwarf galaxy abundances in a qualitatively similar fashion. Specifically, the only difference between the effects of these cutoffs in our inference is the shape of the suppression in dwarf galaxy abundances as a function of $M_{\mathrm{peak}}$, dictated by the astrophysical and WDM cutoffs in Equations 2 and 5, respectively (see Figure 2).
@@
The shape of the WDM SHMF cutoff has been studied in simulations (e.g., Angulo et al. 2013; Lovell et al. 2014; Bose et al. 2016; Stücker et al. 2022), and the shape of the astrophysical cutoff in galaxy abundances has been explored in SAMs (e.g., Benitez-Llambay & Frenk 2020; Kravtsov & Manwadkar 2022; Ahvazi et al. 2023) and hydrodynamic simulations (e.g., Sawala et al. 2016; Fitts et al. 2017; Munshi et al. 2021). Refining predictions for the shapes of both cutoffs and their dependence on halo properties beyond $M_{\mathrm{peak}}$ is an interesting area for future study and will help constrain our model. For example, halo masses evaluated during the epoch of reionization are expected to correlate more strongly with the galaxy occupation fraction than $M_{\mathrm{peak}}$ (e.g., Benitez-Llambay & Frenk 2020), which is typically achieved after reionization. We expect that incorporating such physically-motivated parameterizations will help differentiate astrophysical and DM-induced cutoffs.
@@
In addition to degeneracies with parameters related to our satellite luminosity model, galaxy–halo size connection parameters are generally constrained less precisely than in the scenario without a WDM cutoff. For example, degeneracies between the size amplitude $\mathcal{A}$ or size scatter $\sigma_{\log R}$ and $\mathcal{M}_{50}$ are exacerbated by the larger uncertainty on $M_{\mathrm{hm}}$ in the presence of both astrophysical and DM cutoffs. Physically, degeneracies with the size model arise because our mock observations are surface brightness-limited, such that increasing the sizes of all galaxies at fixed galaxy–halo connection parameters can mimic a cutoff by making these systems too spatially extended to be detectable at a given luminosity (see Nadler et al. 2020b for related discussion).
@@ C.3 Subhalo Mass Function
The posteriors from our one and two-host SHMF forecasts assuming a “weak” galaxy formation cutoff and CDM subhalo abundances are shown in Figures 20 and 21, and the corresponding “strong” galaxy formation cutoff runs are shown in Figures 22 and 23, respectively; the description below applies to both scenarios, which display similar degeneracies. We find that it is particularly challenging to sample the posterior using our MCMC method for SHMF forecasts that combine two hosts; as a result, parameters like $\alpha$ are biased at the $1\sigma$ to $2\sigma$ level in our two-host forecasts (see Figures 21 and 23). However, these biases do not affect the interpretation of degeneracies presented below or the precision of our projected SHMF constraints.
The SHMF amplitude at $M_{\mathrm{peak}}=10^{10}~{}M_{\mathrm{\odot}}$, $\xi_{10}$, does not show noticeable degeneracies with galaxy–halo connection or other SHMF parameters. As described in Section 6, this follows because ($i$) galaxy–halo connection parameters affect satellites that occupy lower-mass halos most strongly, ($ii$) faint satellites are more abundant and thus contribute more to the likelihood, and ($iii$) the effects of varying $\xi_{10}$ cannot be hidden by non-detections, since satellites hosted by the most massive subhalos are always detectable given our observational assumptions.
Several degeneracies appear between $\xi_{8}$, $\xi_{9}$, and the galaxy–halo connection model. Both of these SHMF amplitude parameters are positively correlated with the faint-end luminosity function slope. In our model, $\alpha$ only affects the galaxy–halo connection for $M_{V}>-13~{}\mathrm{mag}$, which roughly corresponds to halos with peak virial masses between $10^{9}~{}M_{\mathrm{\odot}}$ to $10^{10}~{}M_{\mathrm{\odot}}$ (e.g., Nadler et al. 2020b). Since less negative values of $\alpha$ yield fewer faint satellites that occupy such halos, increasing the underlying SHMF amplitude counteracts this effect. In turn, the faint-end slope is measured less precisely in our SHMF forecasts relative to our galaxy formation cutoff and WDM forecasts, and also displays mild degeneracies with galaxy–halo connection parameters like $\sigma_{M}$ when the SHMF is allowed to vary.

### v2

@@ 9 Summary
We have forecasted the galaxy formation and DM constraints that future surveys of dwarf galaxy populations around MW-mass hosts can deliver. We find the following:
@@ Acknowledgements
Our galaxy–halo connection framework is publicly available at github.com/eonadler/subhalo_satellite_connection.
We are grateful to the referee for providing constructive feedback. We thank Keith Bechtol and Alex Drlica-Wagner for comments on the manuscript and Niusha Ahvazi, Rui An, Shany Danieli, Ivan Esteban, and Josh Simon for helpful discussions related to this work.
V.G. acknowledges support from NASA through the Astrophysics Theory Program, award No. 21-ATP21-0135 and from the National Science Foundation (NSF) CAREER grant No. PHY-2239205. The work of LAM was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA. This work received additional support from the U.S. Department of Energy under contract No. DE-AC02-76SF00515 to SLAC National Accelerator Laboratory and from the Kavli Institute for Particle Astrophysics and Cosmology at Stanford University. This work was performed in part at the Aspen Center for Physics, which is supported by NSF grant PHY-2210452.
The computations presented here were conducted through Carnegie’s partnership with the Resnick High Performance Computing Center, a facility supported by the Resnick Sustainability Institute at the California Institute of Technology. This work used data from the Symphony suite of simulations (http://web.stanford.edu/group/gfc/symphony/), which was supported by the Kavli Institute for Particle Astrophysics and Cosmology at Stanford University and SLAC National Accelerator Laboratory, and by the U.S. Department of Energy under contract No. DE-AC02-76SF00515 to SLAC National Accelerator Laboratory.
This research used https://arXiv.org and NASA’s Astrophysics Data System for bibliographic information.
@@ Appendix C Posterior Distributions
Finally, we present the full posteriors from each forecast described in the main text. For all posteriors below, dark (light) contours represent $68\%$ ($95\%$) confidence intervals, and dashed black lines show the input value of each parameter; contours are colored blue (red) for our one-host (two-host) forecasts. Note that $\sigma_{M}$ and $\sigma_{\log R}$ are reported in $\rm{dex}$, $\mathcal{M}_{50}$ is reported as $\log(\mathcal{M}_{50}/M_{\rm{\odot}})$, $\mathcal{A}$ is reported in parsecs, $M_{\mathrm{hm}}$ is reported as $\log(M_{\mathrm{hm}}/M_{\rm{\odot}})$, and the remaining galaxy–halo connection and beyond-CDM parameters are dimensionless. Furthermore, the two-host posteriors below always combine hosts with different bright satellite populations (i.e., classical satellite abundances that differ at a level comparable to the $1\sigma$ host-to-host scatter in our simulation suite).
@@ C.1 Galaxy Formation Cutoff
The posteriors from our one- and two-host “strong” galaxy formation cutoff forecasts are shown in Figures 10 and 11, respectively. The parameters of interest for this forecast (i.e., $\sigma_{M}$, $\mathcal{M}_{50}$, and $\mathcal{S}_{\mathrm{gal}}$) are not strongly correlated with the other galaxy–halo connection parameters; however, there is a prominent degeneracy between $\mathcal{M}_{50}$ and $\sigma_{M}$ (and, to a lesser extent, between $\mathcal{S}_{\mathrm{gal}}$ and $\sigma_{M}$). As described in Section 4, this follows because larger values of $\sigma_{M}$ preferentially cause faint galaxies to up-scatter to observable luminosities, decreasing the average masses of halos inferred to host the faintest observed galaxies. Increasing $\mathcal{M}_{50}$ cuts off galaxy formation more drastically in lower-mass halos, which is not allowed in regions of parameter space where the galaxies they host are crucial to explain the data, resulting in a negatively sloped degeneracy between $\mathcal{M}_{50}$ and $\sigma_{M}$.
Next, the posteriors from our one- and two-host “weak” galaxy formation cutoff forecasts are shown in Figures 12 and 13, respectively. In the one-host case, we only recover an upper limit on $\mathcal{M}_{50}$, and we observe weaker degeneracies between $\mathcal{M}_{50}$ and the remaining galaxy–halo connection parameters. Although uncertainties significantly shrink in the two-host forecast, $\mathcal{M}_{50}$ is still not constrained from below in this case, which is consistent with our discussion in Section 4. Note that the upper limit on $\mathcal{M}_{50}$ in the two-host forecast is not in significant tension with its input value.
@@
Free Parametertexttttt Prior Distributiontext Motivation
@@
WDM half-mode mass $\log(M_{\mathrm{hm}}/M_{\rm{\odot}})\sim\rm{unif}(5,10)$ Brackets current constraints (Gilman et al., 2020a; Banik et al., 2021b)
@@ C.2 Warm Dark Matter
The posteriors from our one- and two-host WDM forecasts assuming a “weak” galaxy formation cutoff and CDM subhalo abundances are shown in Figures 14 and 15, and the corresponding “strong” galaxy formation cutoff runs are shown in Figures 16 and 17, respectively. The posteriors are qualitatively similar in these scenarios, and our description below applies to both.
@@
All of the same degeneracies among galaxy–halo connection parameters noted in Appendix C.1 remain when $M_{\mathrm{hm}}$ is added to the inference. In addition to the significant degeneracies between $M_{\mathrm{hm}}$ and both $\mathcal{M}_{50}$ and $\sigma_{M}$ discussed in Section 5, these posteriors also illustrate a weak degeneracy between $M_{\mathrm{hm}}$ and $\mathcal{S}_{\mathrm{gal}}$ that persists even when two complete satellite populations with different bright satellite populations are used. Figures 16 and 17 clearly illustrate that the $\mathcal{M}_{50}$$\sigma_{M}$ degeneracy is broken when combining two complete satellite populations with differing classical satellite abundances, which in turn yields a much stronger upper limit on $M_{\mathrm{hm}}$ in the two-host case.
@@
Figures 18 and 19, respectively, show the posteriors from our one and two-host WDM forecasts assuming a “strong” galaxy formation cutoff and WDM subhalo abundances with $M_{\mathrm{hm}}=10^{8}~{}M_{\mathrm{\odot}}$ ($m_{\mathrm{WDM}}=4.9~{}\mathrm{keV}$). The degeneracies described in the previous paragraph all remain in this case and often become more prominent. As discussed in Section 5, this follows because, in our model, astrophysical and DM-induced cutoffs affect dwarf galaxy abundances in a qualitatively similar fashion. Specifically, the only difference between the effects of these cutoffs in our inference is the shape of the suppression in dwarf galaxy abundances as a function of $M_{\mathrm{peak}}$, dictated by the astrophysical and WDM cutoffs in Equations 2 and 5, respectively (see Figure 2).
@@
The shape of the WDM SHMF cutoff has been studied in simulations (e.g., Angulo et al. 2013; Lovell et al. 2014; Bose et al. 2016; Stücker et al. 2022), and the shape of the astrophysical cutoff in galaxy abundances has been explored in SAMs (e.g., Benitez-Llambay & Frenk 2020; Kravtsov & Manwadkar 2022; Ahvazi et al. 2024) and hydrodynamic simulations (e.g., Sawala et al. 2016; Fitts et al. 2017; Munshi et al. 2021). Refining predictions for the shapes of both cutoffs and their dependence on halo properties beyond $M_{\mathrm{peak}}$ is an interesting area for future study and will help constrain our model. For example, halo masses evaluated during the epoch of reionization are expected to correlate more strongly with the galaxy occupation fraction than $M_{\mathrm{peak}}$ (e.g., Benitez-Llambay & Frenk 2020), which is typically achieved after reionization. We expect that incorporating such physically motivated parameterizations will help differentiate astrophysical and DM-induced cutoffs.
@@
In addition to degeneracies with parameters related to our satellite luminosity model, galaxy–halo size connection parameters are generally constrained less precisely than in the scenario without a WDM cutoff. For example, degeneracies between the size amplitude $\mathcal{A}$ or size scatter $\sigma_{\log R}$ and $\mathcal{M}_{50}$ are exacerbated by the larger uncertainty on $M_{\mathrm{hm}}$ in the presence of both astrophysical and DM cutoffs. Physically, degeneracies with the size model arise because our mock observations are surface-brightness-limited, such that increasing the sizes of all galaxies at fixed galaxy–halo connection parameters can mimic a cutoff by making these systems too spatially extended to be detectable at a given luminosity (see Nadler et al. 2020b for related discussion).
@@ C.3 Subhalo Mass Function
The posteriors from our one- and two-host SHMF forecasts assuming a “weak” galaxy formation cutoff and CDM subhalo abundances are shown in Figures 20 and 21, and the corresponding “strong” galaxy formation cutoff runs are shown in Figures 22 and 23, respectively; the description below applies to both scenarios, which display similar degeneracies. We find that it is particularly challenging to sample the posterior using our MCMC method for SHMF forecasts that combine two hosts; as a result, parameters like $\alpha$ are biased at the $1\sigma$$2\sigma$ level in our two-host forecasts (see Figures 21 and 23). However, these biases do not affect the interpretation of degeneracies presented below or the precision of our projected SHMF constraints.
The SHMF amplitude at $M_{\mathrm{peak}}=10^{10}~{}M_{\mathrm{\odot}}$, $\xi_{10}$, does not show noticeable degeneracies with galaxy–halo connection or other SHMF parameters. As described in Section 6, this follows because (i) galaxy–halo connection parameters affect satellites that occupy lower-mass halos most strongly, (ii) faint satellites are more abundant and thus contribute more to the likelihood, and (iii) the effects of varying $\xi_{10}$ cannot be hidden by nondetections, since satellites hosted by the most massive subhalos are always detectable given our observational assumptions.
Several degeneracies appear between $\xi_{8}$, $\xi_{9}$, and the galaxy–halo connection model. Both of these SHMF amplitude parameters are positively correlated with the faint-end luminosity function slope. In our model, $\alpha$ only affects the galaxy–halo connection for $M_{V}>-13~{}\mathrm{mag}$, which roughly corresponds to halos with peak virial masses between $10^{9}~{}M_{\mathrm{\odot}}$ and $10^{10}~{}M_{\mathrm{\odot}}$ (e.g., Nadler et al. 2020b). Since less negative values of $\alpha$ yield fewer faint satellites that occupy such halos, increasing the underlying SHMF amplitude counteracts this effect. In turn, the faint-end slope is measured less precisely in our SHMF forecasts relative to our galaxy formation cutoff and WDM forecasts, and also displays mild degeneracies with galaxy–halo connection parameters like $\sigma_{M}$ when the SHMF is allowed to vary.